OpenStudy (anonymous):

7/8 divided by (1/2) to the second power

OpenStudy (anonymous):

show your work

OpenStudy (anonymous):

1/2 squared or the 7/8 /1/2 squared?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

1/2*1/2= 1/4 1/4*7/8=7/32

OpenStudy (anonymous):

so final answer 7/32?

OpenStudy (anonymous):

If ur new, u give medals to whoever answerd it cuz that tells people that its already got a good answer

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

can that be reduced?

OpenStudy (anonymous):

im not new but illg ive u a medal

OpenStudy (anonymous):

ONe sec, thats wrong.

OpenStudy (anonymous):

comeon

OpenStudy (anonymous):

1/4 DIVEDID by 7/8 = 1/4*8/7 =8/28=4/14/2/7! yay!

OpenStudy (anonymous):

SO the final answer is 2/7. There ya go!!!